OpenStudy (anonymous):

7th grade FLVS help @freckles

OpenStudy (anonymous):

Prove that the equation\[a^n + b^n = c^n\]has no positive integral triples (a, b, c) as solutions for \(n \ge 3\) and \(n \in \mathbb N\).

imqwerty (imqwerty):

i saw this one 4 days ago :)

Parth (parthkohli):

pls tell the proof here

OpenStudy (anonymous):

n > 2 it says

OpenStudy (anonymous):

There is not enough space to fit it in this text box. :)

imqwerty (imqwerty):

(B FLT is beyond 7th grade scope i ws trynna get a basic solution lol

Parth (parthkohli):

^ Hah.

OpenStudy (alexandervonhumboldt2):

i like your username

Parth (parthkohli):

ty

OpenStudy (alexandervonhumboldt2):

i wonder why parths dog is so popular

Parth (parthkohli):

because that is me.

OpenStudy (freckles):

i'm having trouble with just the case n=3

Parth (parthkohli):

Ahahaha.

OpenStudy (freckles):

I played with some stuff and have \[a+b=\frac{c^3}{(a+b)^2-3ab} \\ \text{ and since } a \text{ and } b \text{ are assumed to be integers } \\ \text{ then } a+b \text{ is also }\] But I don't if there will be at least one case where (a+b)^2-3ab divides c^3 or not.

OpenStudy (freckles):

I guess I was looking for a proof by contradiction

Parth (parthkohli):

It took three-hundred years to prove this statement ever since it was proposed. Not that easy.

OpenStudy (freckles):

oh it has been proved somewhere?

OpenStudy (freckles):

He's so fancy.

OpenStudy (freckles):

I wish I could be that fancy in math.

Parth (parthkohli):

``` Wiles stumbled upon a revelation, "so indescribably beautiful... so simple and so elegant," ``` I don't find using elliptic curves in number theory elegant. lol But really, what an achievement it was! Great man.

OpenStudy (freckles):

lol

OpenStudy (freckles):

i think i will never graduate from elementary number theory.